What is the magnitude of an electric field that exerts a force of 5.0×10^-4N on a charge of 1.0 uC?

1
Expert's answer
2022-03-20T18:51:18-0400

The magnitude of an electric field

"E=\frac{F}{q}=\rm \frac{5.0\times 10^{-4}\: N}{1.0\times 10^{-6}\: C}=500\: N/C"